OpenStudy (anonymous):

What is the value of x? x+15=8

OpenStudy (anonymous):

ok do u know how simplify expression " and plz no one say the answer"

pooja195 (pooja195):

x+15=8 Start by subtracting both sides by 15 (to cancel out the 15 and to get x alone) x+15-15=8-15 8-15=?

OpenStudy (anonymous):

-7

pooja195 (pooja195):

Good and that would be our answer :-)

OpenStudy (anonymous):

thank youuu